# Communication-Efficient Stochastic Distributed Learning

### v1

@@ III-B Convergence with Variance Reduction
III-C Discussion
III-C1 Choice of step-size
The upper bounds to the step-sizes of LT-ADMM and LT-ADMM-VR ((10) and (11) in Appendix A-A), highlight a dependence on several features of the problem. In particular, the step-size bounds decrease as the smoothness constant $L$ increases, as is usually the case for gradient-based algorithms. Moreover, the bounds are proportional to the network connectivity, represented by the second-largest eigenvalue of $\mathcal{G}$’s Laplacian ($|\tilde{\lambda}_{\max}|$). Thus, less connected graphs (smaller $|\tilde{\lambda}_{\max}|$) result in smaller bounds. Finally, we remark that the step-size bound for LT-ADMM-VR is proportional to $\frac{m_{l}}{m_{u}}=\frac{\min_{i\in\mathcal{V}}m_{i}}{\max_{i\in\mathcal{V}}m_{i}}$, where $m_{i}$ is the number of data points available to agent $i$ (see (1)). This ratio can be viewed as a measure of heterogeneity between the agents. Smaller values of $\frac{m_{l}}{m_{u}}$ highlight higher imbalance in the amount of data available to the agents. The step-size bound thus is smaller for less balanced scenarios.
@@ III-C1 Choice of step-size
III-C2 Convergence rates comparison
As discussed in section I-A, various distributed algorithms with variance reduction have been recently proposed, for example, [12, 14] for strongly convex problems, and [16, 15, 13] for nonconvex problems. Focusing on [16, 15, 13], we notice that their convergence rate is $\mathcal{O}(\frac{1}{K})$, while Theorem 2 shows that LT-ADMM-VR has rate of $\mathcal{O}(\frac{1}{\tau K})$. This shows that employing local training accelerates convergence.
@@ III-C2 Convergence rates comparison
III-C3 Choice of variance reduction mechanism
In variance reduction, we distinguish two main classes of algorithms: those that need to periodically (or randomly) perform a full gradient evaluation (SARAH-type [35]), and those that do not (SAGA-type [25]). In distributed learning, SARAH-type algorithms were proposed in e.g., [16, 14], while SAGA-type algorithms in e.g. [14]. Also the proposed LT-ADMM-VR requires a periodic full gradient evaluation, as the agents re-initialize their local gradient memory at the start of each local training (since they set $r_{i,h,k}^{0}=x_{i,k}$). Clearly, periodically computing a full gradient significantly increases the computational complexity of the algorithm. Thus, one can design a SAGA-type variant of LT-ADMM-VR by removing the gradient memory re-initialization at the start of local training (choosing now $r_{i,h,k}^{0}=r_{i,h,k-1}^{\tau}$). This variant is computationally cheaper and shows promising empirical performance, see the results for LT-ADMM-VR v2 in section IV. However, using the outdated gradient memory leads to a more complex theoretical analysis, which we leave for future work.
III-C4 Uncoordinated parameters
In principle, the agents could use uncoordinated parameters, based on the resources available to them. For example, different agents could employ different local solvers and batch sizes. Alternatively, they could use the same solver but with different step-sizes tailored to the smoothness of the local costs.
IV Numerical Results
@@
where $[x]_{\ell}$ is the $\ell$-th component of $x\in\mathbb{R}^{n}$, and $a_{i,h}\in\mathbb{R}^{n}$ and $b_{i,h}\in\{-1,1\}$ are the pairs of feature vector and label, randomly generated with scikit-learn. We choose a ring graph with $N=10$, and set $n=5$, $m_{i}=100$, $\epsilon=0.01$; the initial conditions are randomly chosen as $x_{i,0}\sim\mathcal{N}(0,100I_{n})$. We use stochastic gradients with a batch of $|\mathcal{B}|=1$. All results are averaged over $100$ Monte Carlo iterations. For the algorithms with local training we select $\tau=5$. We also tune the step-sizes of all algorithms to ensure best performance. Finally, as performance metric we employ $\left\lVert\nabla F(\bar{x}_{k})\right\rVert^{2}$, which is zero if the agents have reached a solution of (2).
We compare LT-ADMM and LT-ADMM-VR with local training algorithms LED [10] and K-GT [9], as well as variance reduction algorithms GT-SARAH [16], and GT-SAGA [14]. We also compare with the alternative version LT-ADMM-VR v2 discussed in section III-C3. We account for the computation time (in unspecified units) of each algorithm when scaling the x-axis. This choice deviates from the common approach of using the iteration count for the x-axis, since we are interested in simulating the real performance of these algorithms. In particular, letting $t_{\mathrm{G}}$ be the time for a component gradient evaluation ($\nabla f_{i,h}$), and $t_{\mathrm{C}}$ the time for a round of communications, Table II reports the computation time incurred by each algorithm over the course of $\tau$ iterations.
TABLE II: Computation time of the algorithms over $\tau$ iterations.
@@
Figure 1 depicts the comparison of the algorithms in Table II in terms of $\left\lVert\nabla F(\bar{x}_{k})\right\rVert^{2}$. The x-axis for each algorithm is scaled to account for the individual computation time, and we plot the results for three choice of the $t_{\mathrm{G}}/t_{\mathrm{C}}$ ratio.
Figure 1: Algorithms’ comparison for different $t_{\mathrm{G}}/t_{\mathrm{C}}$ ratios. The x-axis is in unspecified computation time units to account for the algorithms’ computational complexity.
First of all, we notice that the algorithms without variance reduction (LED, K-GT, LT-ADMM) as expected only converge to a neighborhood of the optimal solution, as stochastic gradients are employed. The asymptotic errors attained by the algorithms are: LED $8.169\times 10^{-4}$, K-GT $7.552\times 10^{-4}$, LT-ADMM $6.327\times 10^{-4}$.
Turning to the algorithms with variance reduction, we see that all of them converge to the optimal solution; however, depending on the value of $t_{\mathrm{G}}/t_{\mathrm{C}}$, their relative speed of convergence changes. When gradient computations are cheaper than communications ($t_{\mathrm{G}}/t_{\mathrm{C}}=0.1$), the proposed algorithm LT-ADMM-VR (and LT-ADMM-VR v2) outperform both GT-SARAH and GT-SAGA, since the latter do not employ local training. This testifies to the benefit of employing local training in scenarios where communications are expensive. As the ratio $t_{\mathrm{G}}/t_{\mathrm{C}}$ increases to $1$ and then $10$, we see how LT-ADMM-VR and GT-SARAH, on the one hand, and LT-ADMM-VR v2 and GT-SAGA, on the other hand, tend to align in terms of performance, as the bulk of the computation time is now due to gradient evaluations, of which the two pairs of algorithms have a similar number (see Table II).
We conclude evaluating the convergence speed of LT-ADMM-VR for different choices of the number of local training epochs $\tau$. Figure 2 reports the computation time (scaled according to Table II) to reach $\left\lVert\nabla F(\bar{x}_{k})\right\rVert^{2}<10^{-9}$.
Figure 2: Computation time needed by LT-ADMM-VR to reach $\left\lVert\nabla F(\bar{x}_{k})\right\rVert^{2}<10^{-9}$ for different numbers of local training epochs $\tau$.
@@
V Concluding Remarks
In this paper, we considered (non)convex distributed learning problems. In particular, to address the challenge of expensive communication, we proposed two communication-efficient algorithms, LT-ADMM and LT-ADMM-VR, that use local training. The algorithms employ SGD and SGD with variance reduction, respectively. We have shown that LT-ADMM converges to a neighborhood of a stationary point, while LT-ADMM-VR converges exactly. We have thoroughly compared our algorithms with the state of the art, both theoretically and in simulations. Future research will focus on analyzing convergence for strongly convex problems, and extending our algorithm design to asynchronous scenarios.
Appendix A Preliminary Analysis
@@
A-A Step-size bounds
@@
$\bar{\gamma}_{\text{sgd}}\coloneqq\min_{i=1,2,\ldots,7}\bar{\gamma}_{i},$ (10)
$\bar{\gamma}_{\text{sarah}}\coloneqq\min_{i=1,8,9,\ldots,17}\bar{\gamma}_{i},$ (11)
@@
$\bar{\gamma}_{1}$ $\coloneqq\min\left\{1,\frac{2}{|\tilde{\lambda}_{\min}|\rho\tau}\right\},\ \bar{\gamma}_{2}\coloneqq\frac{1}{4\sqrt{\tau(\tau-1)(L^{2}+\rho^{2}d_{u}^{2})}},$
$\bar{\gamma}_{3}$ $\coloneqq{\frac{3}{16L\tau}},\quad\bar{\gamma}_{4}\coloneqq\frac{1}{8\tau\sqrt{L^{2}+4L\rho^{2}d_{u}^{2}\tau+4\rho^{2}d_{u}^{2}}},$
$\bar{\gamma}_{5}$ $\coloneqq\frac{|\tilde{\lambda}_{\max}|^{2}\rho^{2}}{4(\beta_{0}\kappa_{1}+\beta_{1})},\bar{\gamma}_{6}\coloneqq\frac{|\tilde{\lambda}_{\max}|^{2}\rho^{2}}{8(\beta_{0}\kappa_{1}+\beta_{1})},$
$\bar{\gamma}_{7}$ $\coloneqq\max\left\{\frac{|\tilde{\lambda}_{\max}|^{2}\rho^{2}}{256\kappa_{4}\tau^{2}N\beta_{0}},\frac{|\tilde{\lambda}_{\max}|^{2}\rho^{2}\tau^{2}}{48L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2}N}\right\}$
$\bar{\gamma}_{8}$ $\coloneqq\frac{m_{l}}{8L^{2}m_{u}},\quad\bar{\gamma}_{9}\coloneqq\min\left\{\frac{1}{2(4L^{2}+2\rho^{2}d_{u}^{2})},\frac{1}{2}\right\}$
$\bar{\gamma}_{10}$ $\coloneqq\frac{\sqrt{m_{l}}}{\tau L\sqrt{384m_{u}}},\quad\bar{\gamma}_{11}\coloneqq\frac{1}{2\sqrt{6\tau(\tau-1)(3L^{2}+\rho^{2}d_{u}^{2})}}$
$\bar{\gamma}_{12}$ $\coloneqq\frac{1}{8\sqrt{2\alpha_{0}\tau^{2}N+(12\alpha_{0}\tau L^{2}+\frac{L^{3}}{N})\tilde{s}_{2}}},\quad\bar{\gamma}_{13}\coloneqq{\frac{3}{16L\tau}}$
$\bar{\gamma}_{14}$ $\coloneqq\frac{|\tilde{\lambda}_{\max}|^{2}\rho^{2}}{4(\tilde{\beta}_{0}\kappa_{1}+\tilde{\beta}_{1}+(96\tau^{2}L^{2}\tilde{\beta}_{0}+2\tilde{\beta}_{3})(\tilde{s}_{0}+\tilde{s}_{1}))}$
$\bar{\gamma}_{15}$ $\coloneqq\frac{|\tilde{\lambda}_{\max}|^{2}\rho^{2}}{8(\tilde{\beta}_{0}\kappa_{1}+\tilde{\beta}_{1}+(96\tau^{2}L^{2}\tilde{\beta}_{0}+2\tilde{\beta}_{3})(\tilde{s}_{0}+\tilde{s}_{1}))}$
@@
$\bar{\gamma}_{16}$ $\coloneqq\frac{24L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2}N}{32\tau^{2}N\tilde{\beta}_{0}+192\tilde{s}_{2}\tau^{2}L^{2}\tilde{\beta}_{0}+4\tilde{s}_{2}\tilde{\beta}_{3}}$
@@
$\bar{\gamma}_{17}$ $\coloneqq\frac{|\tilde{\lambda}_{\max}|^{2}\rho}{48\mu L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2}N}.$
These bounds depend on the following quantities. $d_{u}=\max\{|\mathcal{N}_{i}|\}_{{i\in\mathcal{V}}}$ denotes the maximum agents’ degree. $\widehat{\mathbf{V}}$ is defined in (25). $\tilde{\lambda}_{\min}=\lambda_{\min}(\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}})$ and $\tilde{\lambda}_{\max}=\lambda_{\max}(\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}})$, with $-\tilde{\mathbf{L}}$ being the graph Laplacian (see (15)), and $\widehat{\mathbf{Q}}$ is being defined at the beginning of Appendix A-C. We denote $m_{u}=\max_{i=1,...,N}m_{i}$ and $m_{l}=\min_{i=1,...,N}m_{i}$, where $m_{i}$ is the number of local data points of agent $i$. Additionally, we have the following definitions, used both in the upper bound above and throughout the convergence analysis:
$\beta_{0}$ $\coloneqq 4(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})\|\widehat{\mathbf{V}}^{-1}\|^{2}(L^{2}+\rho^{2}d_{u}^{2})$
$+{12L^{2}}\|\widehat{\mathbf{V}}^{-1}\|^{2}\rho^{2}d_{u}^{2}$
@@
$\beta_{1}$ $\coloneqq(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})\|\widehat{\mathbf{V}}^{-1}\|^{2}\tau\rho d_{u}^{2}{72}/{|\tilde{\lambda}_{\max}|}$
$+L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2}18\rho d_{u}^{2}\tau{216}/{|\tilde{\lambda}_{\max}|}$

### v2

@@ 3.2 Convergence with variance reduction
3.3 Discussion
3.3.1 Choice of step-size
The upper bounds to the step-sizes of LT-ADMM and LT-ADMM-VR ((10) and (11) in Appendix 7.1), highlight a dependence on several features of the problem. In particular, the step-size bounds decrease as the smoothness constant $L$ increases, as is usually the case for gradient-based algorithms. Moreover, the bounds are proportional to the network connectivity, represented by the smallest nonzero eigenvalue of $\mathcal{G}$’s Laplacian (the algebraic connectivity $\lambda_{l}$). Thus, less connected graphs (smaller $\lambda_{l}$) result in smaller bounds. Finally, we remark that the step-size bound for LT-ADMM-VR is proportional to $\frac{m_{l}}{m_{u}}=\frac{\min_{i\in\mathcal{V}}m_{i}}{\max_{i\in\mathcal{V}}m_{i}}$, where $m_{i}$ is the number of data points available to agent $i$ (see (1)). This ratio can be viewed as a measure of heterogeneity between the agents. Smaller values of $\frac{m_{l}}{m_{u}}$ highlight higher imbalance in the amount of data available to the agents. The step-size bound thus is smaller for less balanced scenarios.
@@ 3.3.1 Choice of step-size
3.3.2 Convergence rates
As discussed in section 1.1, various distributed algorithms with variance reduction have been recently proposed, for example, [12, 14] for strongly convex problems, and [16, 15, 13] for nonconvex problems. Focusing on [16, 15, 13], we notice that their convergence rate is $\mathcal{O}(\frac{1}{K})$, while Theorem 2 shows that LT-ADMM-VR has rate of $\mathcal{O}(\frac{1}{\tau K})$. This shows that employing local training accelerates convergence.
@@ 3.3.2 Convergence rates
We also remark that, as shown in (48) and (70) in the Appendix, better network connectivity (corresponding to larger $\lambda_{l}$) leads to smaller upper bounds on the right-hand side of the convergence results. This indicates that stronger network connectivity accelerates the convergence rate.
Finally, the bound in Theorem 1 also highlights a trade-off: a larger $\gamma$ accelerates convergence through the term $\mathcal{O}\!(\tfrac{1}{K\gamma\tau})$, but it also enlarges the steady-state neighborhood via the term $\mathcal{O}\!\left(\gamma\tau\sigma^{2}\right)$. Thus, $\gamma$ must be tuned to balance convergence speed and steady-state precision – and a similar discussion holds for $\tau$. This trade-off is also explored in the numerical results of section 4.2.
3.3.3 Choice of variance reduction mechanism
In variance reduction, we distinguish two main classes of algorithms: those that need to periodically (or randomly) perform a full gradient evaluation (SARAH-type [36]), and those that do not (SAGA-type [25]). In distributed learning, SARAH-type algorithms were proposed in e.g., [16, 14], while SAGA-type algorithms in e.g. [14]. Also the proposed LT-ADMM-VR requires a periodic full gradient evaluation, as the agents re-initialize their local gradient memory at the start of each local training (since they set $r_{i,h,k}^{0}=x_{i,k}$). Clearly, periodically computing a full gradient significantly increases the computational complexity of the algorithm. Thus, one can design a SAGA-type variant of LT-ADMM-VR by removing the gradient memory re-initialization at the start of local training (choosing now $r_{i,h,k}^{0}=r_{i,h,k-1}^{\tau}$). This variant is computationally cheaper and shows promising empirical performance, see the results for LT-ADMM-VR v2 in section 4. However, using the outdated gradient memory leads to a more complex theoretical analysis, which we leave for future work.
3.3.4 Uncoordinated parameters
In principle, the agents could employ uncoordinated parameters, depending on their available resources (e.g., heterogeneous computational capabilities). For instance, different agents could adopt distinct local solvers, numbers of updates (see results in Section 4.2), and batch sizes. Alternatively, they could use the same solver but with step-sizes tailored to the smoothness of their local cost functions.
4 Numerical Results
@@
where $x]_{\ell}$ is the $\ell$-th component of $x\in\mathbb{R}^{n}$, and $a_{i,h}\in\mathbb{R}^{n}$ and $b_{i,h}\in\{-1,1\}$ are the pairs of feature vector and label. As data we use $8\times 8$ gray-scale images of handwritten digits [^2] From https://doi.org/10.24432/C50P49., with pixels normalized to $[0,1]$; we divide the images in the two classes ‘even’ and ‘odd’. We choose a ring graph with $N=10$, and have $n=64$, $m_{i}=180$, $\epsilon=0.01$; the initial conditions are randomly chosen as $x_{i,0}\sim\mathcal{N}(0,100I_{n})$. We use stochastic gradients with a batch of $|\mathcal{B}|=1$. All results are averaged over $10$ Monte Carlo iterations. For the algorithms with local training we select $\tau=2$. We also tune the step-sizes of all algorithms to ensure best performance. Finally, as performance metric we employ $\left\lVert\nabla F(\bar{x}_{k})\right\rVert^{2}$, which is zero if the agents have reached a solution of (2) [^3] We choose this metric as it can be defined for all algorithms considered in the comparison, whereas $\mathcal{D}_{k}$ is defined specifically for Algorithm 1. . The simulations are implemented in Python and run on a Windows laptop with Intel i7-1265U and 16GB of RAM.
4.1 Comparison with the state of the art
We start by comparing LT-ADMM and LT-ADMM-VR with local training algorithms LED [10] and K-GT [9], as well as variance reduction algorithms GT-SARAH [16], and GT-SAGA [14]. We also compare with the alternative version LT-ADMM-VR v2 discussed in section 3.3.3. When evaluating the performance, we account for the computation time of each algorithm, rather than the more commonly used iteration count. In particular, letting $t_{\mathrm{G}}$ be the time for a component gradient evaluation ($\nabla f_{i,h}$), and $t_{\mathrm{C}}$ the time for a round of communications, Table 2 reports the computation time incurred by each algorithm over the course of $\tau$ iterations.
Table 2: Computation time of the algorithms over $\tau$ iterations.
@@ 4.1 Comparison with the state of the art
We start by comparing in Table 3 the algorithms with variance reduction, in terms of the computation time they require to reach $\left\lVert\nabla F(\bar{x}_{k})\right\rVert^{2}<10^{-7}$, that is, to reach a stationary point up to numerical precision.
Table 3: Comparison of computation time for variance-reduced algorithms to reach $\left\lVert\nabla F(\bar{x}_{k})\right\rVert^{2}<10^{-7}$.
Algorithm [Ref.] $t_{\mathrm{G}}/t_{\mathrm{C}}=0.1$ $t_{\mathrm{G}}/t_{\mathrm{C}}=1$ $t_{\mathrm{G}}/t_{\mathrm{C}}=10$
GT-SARAH [16] $7.57\times 10^{5}$ $6.33\times 10^{6}$ $6.20\times 10^{7}$
GT-SAGA [14] $1.55\times 10^{5}$ $2.21\times 10^{5}$ $8.85\times 10^{5}$
LT-ADMM-VR $6.04\times 10^{5}$ $5.76\times 10^{6}$ $5.73\times 10^{7}$
LT-ADMM-VR v2 $3.81\times 10^{4}$ $9.52\times 10^{4}$ $6.66\times 10^{5}$
We see that, depending on the ratio $t_{\mathrm{G}}/t_{\mathrm{C}}$, their relative speed of convergence changes. When gradient computations are cheaper than communications ($t_{\mathrm{G}}/t_{\mathrm{C}}=0.1$), the proposed algorithm LT-ADMM-VR (and LT-ADMM-VR v2) outperform both GT-SARAH and GT-SAGA, since the latter do not employ local training. This testifies to the benefit of employing local training in scenarios where communications are expensive. As the ratio $t_{\mathrm{G}}/t_{\mathrm{C}}$ increases to $1$ and then $10$, we see how LT-ADMM-VR and GT-SARAH, on the one hand, and LT-ADMM-VR v2 and GT-SAGA, on the other hand, tend to align in terms of performance, as the bulk of the computation time is now due to gradient evaluations, of which the two pairs of algorithms have a similar number (see Table 2). Nonetheless, local training still gives an edge to the proposed algorithms.
The remaining algorithms (LT-ADMM, LED, K-GT) do not guarantee exact convergence as they do not employ variance reduction. Thus in Table 4 we report the asymptotic value of $\left\lVert\nabla F(\bar{x}_{k})\right\rVert^{2}$ achieved by the different methods.
Table 4: Comparison of algorithms without variance reduction.
Algorithm [Ref.] $\left\lVert\nabla F(\bar{x}_{K})\right\rVert^{2}$
LED [10] $1.29\times 10^{-3}$
K-GT [9] $2.01\times 10^{-3}$
LT-ADMM $1.07\times 10^{-3}$
The algorithms have close performance, with the proposed LT-ADMM outperforming the state of the art slightly, that is, converging closer to a stationary point.
4.2 Tuning the parameters
In this section we focus on evaluating the impact of the proposed algorithms’ tunable parameters. As discussed in section 3.3.2, the step-size of LT-ADMM regulates both the speed of convergence and how close it converges to a stationary point. In Table 5 then we apply different step-sizes and evaluate both the asymptotic value of $\left\lVert\nabla F(\bar{x}_{k})\right\rVert^{2}$ and the computation time needed for LT-ADMM to reach such value.
Table 5: Performance of LT-ADMM for different $\gamma$.
$\gamma$ $\left\lVert\nabla F(\bar{x}_{K})\right\rVert^{2}$ Computation time
$0.1$ $6.01\times 10^{-5}$ $4.80\times 10^{4}$
11 $5.79\times 10^{-4}$ $3.22\times 10^{4}$
22 $1.07\times 10^{-3}$ $2.27\times 10^{2}$
33 $1.72\times 10^{-3}$ $2.38\times 10^{2}$
44 $2.68\times 10^{-3}$ $7.9\times 10^{1}$
55 $4.43\times 10^{-3}$ $4.10\times 10^{1}$
As expected, a smaller step-size leads to a smaller asymptotic distance from the stationary point, while a larger step-size improves the speed of convergence.
We turn now to LT-ADMM-VR and evaluate its speed of convergence for different numbers of local training epochs $\tau$. Figure 1 reports the computation time to reach $\left\lVert\nabla F(\bar{x}_{k})\right\rVert^{2}<10^{-7}$.
Figure 1: Computation time for LT-ADMM-VR to reach $\left\lVert\nabla F(\bar{x}_{k})\right\rVert^{2}<10^{-7}$ for different numbers of local training epochs $\tau$.
@@ 4.2 Tuning the parameters
Finally, as discussed in section 3.3.4, we can actually choose uncoordinated parameters in LT-ADMM-VR. In Table 6 then we test the use of different $\tau_{i}$, $i\in\{1,\ldots,N\}$, for different agents.
Table 6: Computation time of LT-ADMM-VR with uncoordinated numbers of local training epochs.
$\tau_{i}$ Computation time
22 $\forall i$ $6.04\times 10^{5}$
$\begin{cases}2&i<N/2\\ 5&i\geq N/2\end{cases}$ $5.80\times 10^{5}$
55 $\forall i$ $3.75\times 10^{5}$
In particular, we compare the computation time required to reach $\left\lVert\nabla F(\bar{x}_{k})\right\rVert^{2}<10^{-7}$ in two coordinated scenarios and an uncoordinated scenario where half of the agents are “slow” ($\tau_{i}=2$) and half are “fast” ($\tau_{i}=5$). Interestingly, the algorithm still converges even with uncoordinated parameters, with the presence of “fast” agents improving the performance.
5 Concluding Remarks
In this paper, we considered (non)convex distributed learning problems. In particular, to address the challenge of expensive communication, we proposed two communication-efficient algorithms, LT-ADMM and LT-ADMM-VR, that use local training. The algorithms employ SGD and SGD with variance reduction, respectively. We have shown that LT-ADMM converges to a neighborhood of a stationary point, while LT-ADMM-VR converges exactly. We have thoroughly compared our algorithms with the state of the art, both theoretically and in simulations. Future research will focus on analyzing convergence for strongly convex problems and on extending our algorithmic framework to asynchronous scenarios and to the broader class of composite problems, as in [37].
[…]
6 Proof sketch of the main theorems
6.1 Proof sketch of Theorem 1
Step 1 (Lemma 1)
Reformulate the algorithm into a compact linear dynamical system, in which $\mathbf{h}_{k}$ contains all nonlinearities. Decompose the system into average and deviation components via a projection matrix $\widehat{\mathbf{Q}}$. Use graph connectivity to show that the linear part of the deviation system
$\widehat{\mathbf{d}}_{k+1}=\mathbf{\Delta}\widehat{\mathbf{d}}_{k}-\widehat{\mathbf{h}}_{k}$
is stable ($\|\mathbf{\Delta}\|=1-{\lambda_{l}\rho\tau\beta}/{2}<1$) when $\beta<\frac{2}{\tau\lambda_{u}\rho}$ is satisfied.
Step 2 (Lemmas 8.48.6)
Bound the error from local training steps: the deviation of local states $\boldsymbol{\Phi}_{k}^{t}$ from the global average $\mathbf{\bar{X}}_{k}$ as in (29).
$\mathbb{E}\left[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}\right]$ $\leq\left(\frac{72\beta\tau^{2}}{\lambda_{l}\rho}+144\tau^{3}\beta^{2}\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+4N\tau^{2}\gamma^{2}\sigma^{2}$
$+16\tau^{3}\gamma^{2}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}].$
Incorporate this bound into the perturbation term $\widehat{\mathbf{h}}_{k}$ to derive a recursive inequality for the deviation system as in (37),
$\mathbb{E}[\|\widehat{\mathbf{d}}_{k+1}\|^{2}]$
$\leq(\delta+\frac{c_{0}}{1-\delta})\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+\frac{c_{1}}{1-\delta}\mathbb{E}[\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}]$
$+\frac{c_{2}}{1-\delta}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}]+\frac{c_{3}}{1-\delta}\sigma^{2},$
where sufficiently small $\gamma$ ensuring stability.
Step 3 (Theorem 1)
Apply the smoothness inequality to the averaged iterate
$\bar{x}_{k+1}-x^{*}=\bar{x}_{k}-x^{*}-\frac{\gamma}{N}\sum_{t=0}^{\tau-1}\sum_{i=1}^{N}g_{i}(\phi_{i,k}^{t}),$
show a descent in the global objective $F\left(\bar{x}_{k}\right)$.
$\mathbb{E}[F\left(\bar{x}_{k+1}\right)]\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\frac{\gamma\tau}{2}\mathbb{E}[\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}]$
$-\frac{\gamma}{2}(1-2\gamma\tau L)\sum_{t}\mathbb{E}[\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\frac{\gamma L^{2}}{2N}\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]+{\gamma^{2}\tau^{2}L\sigma^{2}}.$
Sum over iterations, pick an appropriate step-size, and use the bounds obtained in Lemmas 8.48.6, leading to the convergence result.
6.2 Proof sketch of Theorem 2
The proof of Theorem 2 follows a structure similar to that of Theorem 1. The key distinction lies in the treatment of the gradient variance. We bound the gradient variance with $t_{k}$,
$\mathbb{E}[\sum_{i}\sum_{t}\|g_{i}\left(\phi_{i,k}^{t}\right)-\nabla f_{i}(\phi_{i,k}^{t})\|^{2}]\leq 2L^{2}\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+2L^{2}\mathbb{E}[{t}_{k}],$
using $\mathbb{E}[\|a-\mathbb{E}[a]\|^{2}]\leq\mathbb{E}[\|a\|^{2}]$ with $a=\nabla f_{i,h}(\phi_{i,k}^{t})-\nabla f_{i,h}(r_{i,h,k}^{t})$. And show that $t_{k}$ can be bound by the deviation bound $\|\widehat{\mathbf{d}}_{k}\|$ and the global gradient at the average state $\nabla F(\bar{x}_{k})$ as in Lemma 9.10. And we further bound the deviation bound $\|\widehat{\mathbf{d}}_{k}\|$ with $t_{k}$ as in Lemma 9.12. As a result, the final convergence expression no longer depends on a constant $\sigma$, and we can have exact convergence.
7 Preliminary Analysis
@@
7.1 Step-size bounds
@@
$\bar{\gamma}_{\text{sgd}}\coloneqq\min_{i=1,2,\ldots,6}\bar{\gamma}_{i},$ (10)
$\bar{\gamma}_{\text{vr}}\coloneqq\min_{i=1,7,8,\ldots,15}\bar{\gamma}_{i},$ (11)
@@
$\bar{\gamma}_{1}$ $\coloneqq\min\left\{1,\frac{1}{L\tau 2\sqrt{2}}\right\},\ \bar{\gamma}_{2}\coloneqq\frac{\sqrt{3}}{8L\tau},$
@@
$\bar{\gamma}_{3}$ $\coloneqq{\frac{3}{8L\tau}},\quad\bar{\gamma}_{4}\coloneqq\frac{\lambda_{l}}{\lambda_{u}L\sqrt{16(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})\tau\|\widehat{\mathbf{V}}^{-1}\|^{2}\beta_{0}}},$
@@
$\bar{\gamma}_{5}$ $\coloneqq\frac{\sqrt{\lambda_{l}}}{4\tau\sqrt{\lambda_{u}L}\sqrt[4]{c_{4}(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})2\|\widehat{\mathbf{V}}^{-1}\|^{2}}},$
@@
$\bar{\gamma}_{6}$ $\coloneqq\frac{\sqrt{\lambda_{l}\beta}}{2\sqrt{\tau\lambda_{u}L}\sqrt[4]{c_{4}6N\|\widehat{\mathbf{V}}^{-1}\|^{2}}},$
$\bar{\gamma}_{7}$ $\coloneqq\frac{1}{4\tau L\sqrt{3}},\bar{\gamma}_{8}\coloneqq\sqrt{\frac{m_{l}}{32L^{2}m_{u}}},\bar{\gamma}_{9}\coloneqq\sqrt{\frac{1}{12L^{2}}},$
$\bar{\gamma}_{10}$ $\coloneqq\sqrt{\frac{m_{l}}{512m_{u}\tau^{3}L^{2}}},\bar{\gamma}_{11}\coloneqq\sqrt{\frac{3\tau}{8\kappa_{3}}},\bar{\gamma}_{12}\coloneqq\frac{3}{8\tau L},$